OpenStudy (anonymous):

2 times the cubed root of 7b-1 equals 4

OpenStudy (jack1):

\[2\times \sqrt[3]{(7b-1)} = 4\]

OpenStudy (jack1):

this??

OpenStudy (jack1):

@randomguy111

OpenStudy (anonymous):

yep!

OpenStudy (jack1):

k, assuming it it... just remember to always balance the 2 sides of your equation so LHS /2 = RHS /2 2 x cubed root of (7b-1) = 4 2 x cubed root of (7b-1) /2 = 4 /2 cubed root of (7b-1) = 2 now cube both sides (^3), so cubed root of (7b-1) = 2 cubed root of (7b-1) ^3 = 2 ^3 (7b-1) = 8 can u take it from here?

OpenStudy (anonymous):

So b would equal 9/7 ?

OpenStudy (jack1):

that what i got, yep

OpenStudy (anonymous):

Awesome! Thank you so much! Hey I might need more of your help lol

OpenStudy (jack1):

sure, just gonna finish helping constance first

OpenStudy (anonymous):

Okay just let me know :)